Find answers to the following. Write and indicate how you solved them.
 Is 9 / 16 equal to 5/9

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

9 / 16, 5 / 9
Convert into like fractions
9 / 16 = (9 / 16) × (9 / 9)
= 81 / 144
5 / 9 = (5 / 9) × (16 / 16)
= 80 / 144
∴81 / 144 ≠ 80 / 144
Hence, 9 / 16 is not equal to 5 / 9

​Thanks